Two 80-0 resistors are connected in series with 0.15 A flowing through them?

Find a. Total Resistance (Rt)

b. Voltage drop (Vt) across resistors

c. Current through the circuit



1
Expert's answer
2021-04-26T09:49:03-0400

To be given in question

R1=R2 "=80\ ohm"

"I=0.15amp"

To be asked in question

(a)total resistance

Rt=?

(b) voltage drop (v)=?

(c) current through the circuit=?

We know that


(a) series combination

"R_{t}=R_{1}+R_{2}"

put value

Rt=160 ohm

(b) voltage drope

"V_{1} =IR_{1}"

"V_{1}=0.15\times80=12Volt"

"V_{2}=IR_{2}"

"V_{2}=0.15\times80=12 Volt"

"V=V_{1}+V_{2}"

"V=12+12=24Volt"

(c) current through the circuit

"I=\frac{V}{R}"

"I=\frac{24}{160} =0.15 amp"